As a first year Algebra student, I am doing a little extra study and have run into a problem I can't solve.
Solve: [2x-5y] = [1] for y
[ x+3y] [0]
I tried to line them up - one under the other, but once I hit the submit button, it throws them out of alignment [x+3y]is directly under [2x-5y] and [0] is directly under [1]. Thank you.

Answer by malakumar_kos@yahoo.com(315)


Solve: [2x-5y] = [1] for y
[ x+3y] [0]
2x-5y = 1...........(1)
x+3y = 0...........(2)
multiply (2) by 2 and subtract (1) and (3)
2(x+3y) = 2(0)
2x+6y = 0..........(3)
2x-5y = 1
2x+6y = 0
- -
------------
0 -11y = 1
------------
y = -1/11
substituting for y in (2) we get x+3y = 0
x = -3y = -3(-1/11) = 3/11
the solution for the eq'ns is x= 3/11 and y = -1/11